OpenStudy (anonymous):

how do you convert the function f(x)=15^x into a logarithmic one?

OpenStudy (kmeis002):

Do you mean, how do you solve for x?

OpenStudy (xapproachesinfinity):

think about this \[a^b=n \Longrightarrow \log_b(n)=a\]

OpenStudy (anonymous):

so then you would get log(base)x(f(x))=15?

OpenStudy (xapproachesinfinity):

this one for natural log \[\ln x=y \Longrightarrow e^y=x\]

OpenStudy (xapproachesinfinity):

no use the equation \[y=15^x\]

OpenStudy (xapproachesinfinity):

eh it really doen't matter what you choose as a base and exponent

OpenStudy (xapproachesinfinity):

oh i made a mistake with the base lol

OpenStudy (solomonzelman):

yes, that is what I was about to post...

OpenStudy (solomonzelman):

I thought \(\large\color{black}{ \displaystyle a^b=n }\) \(\large\color{black}{ \displaystyle \log_aa^b=log_an }\) \(\large\color{black}{ \displaystyle b=log_an }\)

OpenStudy (solomonzelman):

forgot the backwards slash next to logs, sorry for my bad latex.

OpenStudy (xapproachesinfinity):

again i wrote the same thing i must of drunk

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a^b=n }\) \(\large\color{black}{ \displaystyle \log_a(a^b)=\log_a(n) }\) \(\large\color{black}{ \displaystyle b\log_a(a)=\log_a(n) }\) \(\large\color{black}{ \displaystyle b=\log_a(n) }\)

OpenStudy (solomonzelman):

It happens to me all the time, and I would bet more than to you.

OpenStudy (xapproachesinfinity):

yep it should be \[a^b=n \Longrightarrow \log_a(n)=b\]

OpenStudy (xapproachesinfinity):

does not happen to me usually lol just my brain is burned today hahah

OpenStudy (xapproachesinfinity):

so if we have \[y=15^x \Longrightarrow \log_{15}(y)=x\] just switch x and y and you get \[y=\log_{15}x\]

OpenStudy (xapproachesinfinity):

that is the function you are looking for